OpenStudy (anonymous):

translate the following statement to equations let x represent the number in each case twice the sum of a number and 6 is 18

OpenStudy (anonymous):

2x+6=18?

OpenStudy (anonymous):

2x+ 6 = 18 x = 6

OpenStudy (saifoo.khan):

2x+6 = 18 2x = 12 x = 6